# On Utility Maximization under Multivariate Fake Stationary Affine Volterra Models

### 06 Apr 2026 v2

@@ Abstract
This paper is concerned with Merton’s portfolio optimization problem in a Volterra stochastic environment described by a multivariate fake stationary Volterra–Heston model. Due to the non-Markovianity and non-semimartingality of the underlying processes, the classical stochastic control approach cannot be directly applied in this setting. Instead, the problem is tackled using a stochastic factor solution to a Riccati backward stochastic differential equation (BSDE). Our approach is inspired by the martingale optimality principle combined with a suitable verification argument. The resulting optimal strategies for Merton’s problems are derived in semi-closed form depending on the solutions to time-dependent multivariate Riccati-Volterra equations. Numerical results on a two dimensional fake stationary rough Heston model illustrate the impact of stationary rough volatilities on the optimal Merton strategies.
@@ 1 Introduction
Motivated by several important empirical stylized facts about real financial markets such as choice among multiple assets, rough volatility behavior, correlations across stocks or assets and leverage effects (i.e., correlation between a stock and its volatility), multivariate rough volatility models have recently been developed ; see, e.g., Abi Jaber et al. (2019); Tomas and Rosenbaum (2021). In Aichinger and Desmettre (2021), the authors analyze the Merton portfolio problems with power utility for a class of multivariate Volterra Heston models that features both inter-asset correlations and correlation between a stock and its volatility. In the present paper, we solve the Merton portfolio problem for investors with power, exponential and logarithmic utility functions, within the class of the so-called fake stationary multivariate affine Volterra models.
@@
    We introduce a class of multivariate fake stationary affine Volterra stochastic volatility models that capture key stylized features of financial markets, including heterogeneous roughness across assets, possibly stochastic inter-asset correlations, and leverage effects namely, dependence between asset returns and their respective volatilities while maintaining a consistent modeling framework across time scales, from short to long maturities.
@@
    This model preserve analytical tractability, thereby enabling explicit characterization of the optimal investment strategy for the Merton’s problem, despite the intrinsic challenges posed by multivariate non-Markovian dynamics.
Organization of the Work. The outline of the paper is as follows: Section 2 gives an overview of the model which is needed throughout the paper: We introduce the multi asset financial market, where volatility is modeled by a multivariate class of fake stationary Volterra square root process, and we state the optimization problem. For such a market model we consider in section 3 two different approaches to solve the Merton portfolio problem. More precisely, in section 3 we investigate the classical problem of maximizing the expected utility of terminal wealth in a multi-asset fake stationary Volterra–Heston volatility market, for both power, exponential and logarithmic utility preferences. We first adapt the martingale distortion transformation used in Fouque and Hu (2019); Han and Wong (2020) to the (degenerate) multivariate case. However, as it is pointed out in Abi Jaber et al. (2021); Aichinger and Desmettre (2021), this only works if the correlation structure is highly degenerate. Inspired by the techniques used in Bäuerle and Li (2013); Aichinger and Desmettre (2021), we then provide a solution for the Merton portfolio optimization problem for a more general correlation structure using a verification argument. In Section 4, we demonstrate the practical implications of our findings through numerical experiments based on a two-dimensional fake stationary rough Heston model. Finally, Section 5 is devoted to the proofs of the main results.
@@ 2.2 Formulation of the stochastic Market model
driven by a $d$-dimensional Brownian motion $B$, with a $d\times d$-matrix valued continuous stochastic volatility process $\sigma$ whose dynamics is driven by (2.2) and a $\mathbb{R}^{d}$-valued continuous stochastic process $\lambda$, called market price of risk. Here $\mathbb{1}_{d}$} denotes the vector in $\mathbb{R}^{d}$ with all components equal to $1$ and the correlation structure of $W$ with $B$ is given by
@@
From now on, we set $g_{0}(t):=\varphi(t)V_{0}+\int_{0}^{t}K(t-s)\mu(s)ds$ and $Z_{t}:=\int_{0}^{t}DV_{s}ds+\nu\varsigma(s)\sqrt{\operatorname{diag}(V_{s})}dW_{s},$ for all $t\geq 0$ so that Equation (2.2) reads
@@
One notes that for each, $s\leq T$, $(g_{t}(s))_{t\leq s}$ is the adjusted forward process
$g_{t}(s)=\;\mathbb{E}\Big[V_{s}-\int_{t}^{s}K(s-u)DV_{u}du{\ \Big|\ }{\mathcal{F}}_{t}\Big].\vskip-2.84544pt$ (2.20)
@@ 3 Merton’s portfolio problem: Utility maximisation
    A utility function is a strictly increasing and strictly concave function $U:(0,\infty)\to\mathbb{R}\cup\{-\infty\},$ which is continuously differentiable on $(0,\infty)$.
@@
$\mathcal{V}(x_{0},V_{0}):=\sup_{\alpha(\cdot)\in\mathcal{A}}\mathbb{E}\!\left[U\!\left(X_{T}^{\alpha}\right)\right],\;\text{given}\;x_{0}\;\text{and}\;V_{0}.$ (3.21)
with $X^{\alpha}$ the wealth controlled by $\alpha\in\mathcal{A}$, starting from $x_{0}$ at time 0 and $\mathcal{A}$ the subset of controls $\alpha\in\mathcal{A}$ such that the family $\left\{U\!\left(X_{\tau}^{\alpha}\right):\tau\in[0,T]\right\}$ is uniformly integrable. The preferences of the agent are thus described by the utility function $U$. Mathematically, we aim at identifying the optimal value associated with the utility maximization problem defined over a finite horizon $T$ by
$\mathcal{V}_{t}(x,v):=\operatorname*{ess\,sup}_{\alpha\in\mathcal{A}_{t}}\mathbb{E}_{x,v}\!\Big[U\!\Big(X_{T}^{t,x,v,\alpha}\Big)\mid\mathcal{F}_{t}\Big],\qquad x>0,v>0,\;t\in[0,T].$ (3.22)
and the optimal strategy $\alpha^{\ast}$, given the preference described by the utility function $U(\cdot)$. This is a standard stochastic control problem. The set $\mathcal{A}_{t}$ is the class of all admissible strategies from t with zero being an absorbing state for $X_{t}^{\alpha}$ (bankruptcy). Here $\mathbb{E}_{x,v}$ denotes expectation under the conditional distribution given $X_{t}=x$ and $V_{t}=v$, and where the supremum is taken over all admissible portfolio strategies.
The $\mathbb{F}$-adapted process $\{\mathcal{V}_{t},\,0\leq t\leq T\}$ is a supermartingale. Moreover, there exists an optimal control $\alpha^{*}\in\mathcal{A}$ for (3.21) if and only if the martingale property holds, that is, the process $\{\mathcal{V}^{*}_{t},\,0\leq t\leq T\}$ for $\alpha^{*}$ is a martingale. The idea is that an admissible control (or startegy) is optimal333optimal strategy: No expected gain from deviation, this implies martingality if the associated value process is a martingale and for any other admissible control, it is a supermartingale. That is the classic martingale optimality principle, see, e.g., Hu et al. (2005), Pham (2009, Section 6.6.1) or Jeanblanc et al. (2012).
@@
As stated before, Problem (3.21) seen as an optimization problem with state process $X^{\alpha}$ is non-Markovian and the standard stochastic control approach cannot be applied. Heuristically speaking, the non-Markovian and non-semimartingale characteristics of the multivariate fake stationary Volterra Heston model (2.16)-(2.2) are overcome in the degenerate correlation case by applying the martingale optimality principle and constructing an ansatz, which is inspired by the martingale distortion transformation and the exponential-affine representation (see e.g. (Gnabeyeu, 2026, Theorem A.4.)) in term of an auxiliary process denoted $\Gamma\equiv\Gamma(V_{t})$ (more precisely $\Gamma(g_{t})$). In the general correlation case, we provide a verification argument avoiding restrictions on the correlation structure linked to the martingale distortion transformation.
In fact, Fouque and Hu (2019) showed that if the Sharpe-ratio $\lambda$ in (2.14) is bounded and has bounded derivative, then the value process $\mathcal{V}_{t}$ can be expressed as $\mathcal{V}_{t}(x,v)=J_{t}(X_{t}^{\alpha}=x,V_{t}=v)$, where
$J_{t}(X_{t}^{\alpha},V_{t}):=F(t,X_{t}^{\alpha})\Gamma(V_{t})$ (3.23)
@@ 3. $(c)$
    $\alpha(\cdot)$ is $\mathbb{F}$-adapted and $\int^{T}_{0}\left|\alpha_{s}\right|^{2}ds<\infty$, $\mathbb{P}$-a.s.;
@@ 3.1.2 The general correlation case: A verification argument
We will directly show that the value attained by this portfolio strategy is $\frac{x_{0}^{\gamma}}{\gamma}\Gamma_{0}$, and that no other admissible portfolio strategy can yield a larger value, through the following verification result:
@@ 1. $(a)$
    $\alpha(\cdot)$ is $\mathbb{F}$-adapted and $\mathbb{E}[\int^{T}_{0}\left|\alpha_{s}\right|^{2}ds]<\infty$;
@@ 3.2.2 The general correlation case: A verification argument
First, note that if $K_{i}$ satisfies the Assumption 2.1 for $i=1,\ldots,d$. As $1-\rho_{i}^{2}\geq 0$, then (Gnabeyeu, 2026, Theorem A.2.) provides the existence of a unique global continuous solution on $[0,T]$ to (3.67)– (3.68). More details are given in the Remark on Proposition 3.11. In particular, when $\rho_{1}=\ldots=\rho_{d}=\rho$, the function $\psi$ coincides with the unique global solution to the Riccati–Volterra equations (3.59)–(3.60). Consequently, to avoid restrictions on the correlation structure, we still use a verification arguments in the spirit of Bäuerle and Li (2013) to solve the optimization problem, thus extending the results obtained in the preveous section to the more general correlation structure.
@@
For the sake of brevity, yet without compromising self-containment, we present a sketch of the proof.
$G(x_{0},V_{0}):=-\frac{\Gamma_{0}}{\gamma}\exp\big(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\big)=-\frac{1}{\gamma}\exp\big(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}+\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\big)$
we have
  1. 1.
    $\mathbb{E}_{x_{0},V_{0}}{\big[-\frac{1}{\gamma}\exp\big(-\gamma X_{T}^{\alpha^{*}}\big)\big]}=G(x_{0},V_{0})$ for $\alpha_{t}^{*}=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\left(\lambda_{t}+\Sigma\Lambda_{t}\right),\;0\leq t\leq T$,
  2. 2.
    $\mathbb{E}_{x_{0},V_{0}}{\big[-\frac{1}{\gamma}\exp\big(-\gamma X_{T}^{\alpha}\big)\big]}\leq G(x_{0},V_{0})$ for every other admissible strategy $\alpha\in\mathcal{A}$.
In fact, conditions $(1)$ and $(2)$ ensure that G is the value function of the problem (3.57) and $\alpha^{*}$ is the optimal portfolio strategy.

### 07 May 2026 v3

@@ Abstract
This paper is concerned with Merton’s portfolio optimization problem in a Volterra stochastic environment described by a multivariate fake stationary Volterra–Heston model. Due to the non-Markovianity and non-semimartingality of the underlying processes, the classical stochastic control approach cannot be directly applied in this setting. Instead, the problem is tackled using a stochastic factor solution to a Riccati backward stochastic differential equation (BSDE). Our approach is inspired by the martingale optimality principle combined with a suitable verification argument. The resulting optimal strategies for Merton’s problems are derived in semi-closed form depending on the solutions to time-dependent multivariate Riccati-Volterra equations, while the optimal value is expressed using the solution to this original Riccati BSDE. Numerical results on a two dimensional fake stationary rough Heston model illustrate the impact of stationary rough volatilities on the optimal Merton strategies.
@@ 1 Introduction
Motivated by several important empirical stylized facts about real financial markets such as choice among multiple assets, rough volatility behavior, correlations across stocks or assets and leverage effects (i.e., correlation between a stock and its volatility), multivariate rough volatility models have recently been developed ; see, e.g., Abi Jaber et al. (2019); Tomas and Rosenbaum (2021). In Aichinger and Desmettre (2021), the authors analyze the Merton portfolio problems with power utility for a class of multivariate Volterra Heston models that features correlation between a stock and its volatility. In the present paper, we solve the Merton portfolio problem for investors with power, exponential and logarithmic utility functions, within the class of the so-called fake stationary multivariate affine Volterra models.
@@
    We introduce a class of multivariate fake stationary affine Volterra stochastic volatility models that capture key stylized features of financial markets, including heterogeneous roughness across assets, and leverage effects namely, dependence between asset returns and their respective volatilities while maintaining a consistent modeling framework across time scales, from short to long maturities.
@@
    This model preserve analytical tractability, thereby enabling explicit characterization of the optimal investment strategy for the Merton’s problem, despite the intrinsic challenges posed by multivariate non-Markovian dynamics. The optimal value is characterized by the solution of an associated Riccati BSDE (Riccati BSDE) whose generator is quadratic in the Brownian stochastic integrand (denoted by $\Lambda$).
  • (iii)
    It further extends the models of Abi Jaber et al. (2019); Tomas and Rosenbaum (2021) to an inhomogeneous setting, following Gnabeyeu et al. (2025, 2026). Within the framework of Volterra Heston models, our results extend the one-dimensional exponential utility case of Han and Wong (2020) to the multidimensional setting, while encompassing the power utility results of Han and Wong (2020); Aichinger and Desmettre (2021) as special cases. In addition, we incorporate the logarithmic utility case.
Organization of the Work. The outline of the paper is as follows: Section 2 gives an overview of the model which is needed throughout the paper: We introduce the multi asset financial market, where volatility is modeled by a multivariate class of fake stationary Volterra square root process, and we state the optimization problem. For such a market model we consider in section 3 two different approaches to solve the Merton portfolio problem. More precisely, in section 3 we investigate the classical problem of maximizing the expected utility of terminal wealth in a multi-asset fake stationary Volterra–Heston volatility market, for each of the power, exponential and logarithmic utility preferences. We first adapt the martingale distortion transformation used in Fouque and Hu (2019); Han and Wong (2020) to the (degenerate) multivariate case. However, as it is pointed out in Abi Jaber et al. (2021); Aichinger and Desmettre (2021), this only works if the correlation structure is highly degenerate. Inspired by the techniques used in Hu et al. (2005); Bäuerle and Li (2013); Aichinger and Desmettre (2021), we then provide a solution for the Merton portfolio optimization problem for a more general correlation structure using a verification argument. In Section 4, we demonstrate the practical implications of our findings through numerical experiments based on a two-dimensional fake stationary rough Heston model. Finally, Section 5 is devoted to the proofs of the main results.
@@
$\bullet$ Let $M$ denote a continuous semimartingale. The stochastic exponential $\mathcal{E}(M)$ is given by
$\mathcal{E}(M)_{t}=\exp\left(M_{t}-\tfrac{1}{2}\langle M\rangle_{t}\right),\quad t\in[0,T],$
where $\langle M\rangle$ denotes the quadratic variation of $M$.
@@ 2.2 Formulation of the stochastic Market model
driven by a $d$-dimensional Brownian motion $B$, with a $d\times d$-matrix valued continuous stochastic volatility process $\sigma$ whose dynamics is driven by (2.2) and a $\mathbb{R}^{d}$-valued continuous stochastic process $\lambda$, called market price of risk. Here $\mathbb{1}_{d}$} denotes the vector in $\mathbb{R}^{d}$ with all components equal to $1$ and the correlation structure of $W$ with $B$, denoted $\Sigma:=[\Sigma_{1},\cdots,\Sigma_{d}]$ is given by
@@
From now on, we set $g_{0}(t):=\varphi(t)V_{0}+\int_{0}^{t}K(t-s)\mu(s)ds$ the initial input curve and $Z_{t}:=\int_{0}^{t}DV_{s}ds+\nu\varsigma(s)\sqrt{\operatorname{diag}(V_{s})}dW_{s},$ for all $t\geq 0$ so that Equation (2.2) reads $\mathbb{P}\otimes dt$-a.e.
@@
One notes that for each, $0\leq s\leq T$, $(g_{t}(s))_{t\leq s}$ is the adjusted forward process
$g_{t}(s)=\;\mathbb{E}^{\mathbb{P}}\Big[V_{s}-\int_{t}^{s}K(s-u)DV_{u}du{\ \Big|\ }{\mathcal{F}}_{t}\Big],\quad\mathbb{P}\text{- a.s., for a.e.}\;s>t.\vskip-2.84544pt$ (2.20)
@@
[…]
$\widetilde{g}_{t}(s):=\mathbb{E}^{\widetilde{\mathbb{P}}}\Big[V_{s}-\int_{t}^{s}K(s-u)H_{u}V_{u}du|{\mathcal{F}}_{t}\Big],\quad\widetilde{\mathbb{P}}\text{- a.s., for a.e.}\;s>t.\vskip-2.84544pt$ (2.21)
where the dynamics of $V$ under $\widetilde{\mathbb{P}}$ reads
$V_{t}=g_{0}(t)+\int_{0}^{t}K(t-s)d\widetilde{Z}_{s},\quad\text{with}\quad\widetilde{Z}_{t}=\int_{0}^{t}H_{s}V_{s}ds+\nu\varsigma(s)\sqrt{\operatorname{diag}(V_{s})}d\widetilde{W}_{s}.$ (2.22)
@@ 3 Merton’s portfolio problem: Utility maximisation
    A utility function is a strictly increasing and strictly concave function $U:E\subset\mathbb{R}\to\mathbb{R}\cup\{-\infty\},$ which is continuously differentiable on $E$.
@@
$\mathcal{V}(x_{0},V_{0}):=\sup_{\alpha(\cdot)\in\mathcal{A}}\mathbb{E}_{x_{0},V_{0}}\!\left[U\!\left(X_{T}^{\alpha}\right)\right],\;\text{given}\;x_{0}\;\text{and}\;V_{0}.$ (3.23)
with $X^{\alpha}$ the wealth controlled by $\alpha\in\mathcal{A}$, starting from $x_{0}$ at time 0 and $\mathcal{A}$ the subset of controls $\alpha\in\mathcal{A}$ such that the family $\left\{U\!\left(X_{\tau}^{\alpha}\right):\tau\in[0,T]\right\}$ is uniformly integrable. The preferences of the agent are thus described by the utility function $U$. Here $\mathbb{E}_{x_{0},V_{0}}$ denotes expectation under the conditional distribution given $X_{0}=x_{0}$ and $V_{0}$, and the supremum is taken over all admissible portfolio strategies. Mathematically, denoting by $\mathcal{S}_{t}:=\big(X_{t}^{\alpha},g_{t}(\cdot)\big)$ the state variable, we aim at identifying the optimal value associated with the utility maximization problem defined over a finite horizon $T$ by
$\mathcal{V}_{t}(\mathcal{S}_{t}):=\operatorname*{ess\,sup}_{\alpha\in\mathcal{A}_{t}}\mathbb{E}\!\,\big[U\!\big(X_{T}^{\alpha}\big)\mid\mathcal{S}_{t}\big],\qquad t\in[0,T].\vskip-2.84544pt$ (3.24)
and the optimal strategy $\alpha^{\ast}$, given the preference described by the utility function $U(\cdot)$. This is a standard stochastic control problem. The set $\mathcal{A}_{t}$ is the class of all admissible strategies from t with zero being an absorbing state for $X_{t}^{\alpha}$ (bankruptcy).
The $\mathbb{F}$-adapted process $\{\mathcal{V}_{t},\,0\leq t\leq T\}$ is a supermartingale. Moreover, there exists an optimal control $\alpha^{*}\in\mathcal{A}$ for (3.23) if and only if the martingale property holds, that is, the process $\{\mathcal{V}^{*}_{t},\,0\leq t\leq T\}$ for $\alpha^{*}$ is a martingale. The idea is that an admissible control (or startegy) is optimal if the associated value process is a martingale and for any other admissible control, it is a supermartingale. That is the classic martingale optimality principle, see, e.g., Hu et al. (2005), Pham (2009, Section 6.6.1) or Jeanblanc et al. (2012).
@@
As stated before, Problem (3.23) seen as an optimization problem with state process $X^{\alpha}$ is non-Markovian and the standard stochastic control approach cannot be applied. Heuristically speaking, the non-Markovian and non-semimartingale characteristics of the multivariate fake stationary Volterra Heston model (2.16)-(2.2) are overcome in the degenerate correlation case by applying the martingale optimality principle and constructing an ansatz, which is inspired by the martingale distortion transformation and the exponential-affine representation (see e.g. (Gnabeyeu, 2026, Theorem A.4.)) in term of an auxiliary process denoted $\Gamma\equiv\Gamma\big(g_{t}(\cdot)\big)$. In the general correlation case, we provide a verification argument avoiding restrictions on the correlation structure linked to the martingale distortion transformation.
In fact, Fouque and Hu (2019) showed that if the Sharpe-ratio $\lambda$ in (2.14) is bounded and has bounded derivative, then the value process $\mathcal{V}_{t}$ can be expressed as $\mathcal{V}_{t}=J_{t}(X_{t}^{\alpha},g_{t}(\cdot)\big))$, where
$J_{t}(X_{t}^{\alpha},g_{t}(\cdot)\big)):=F(t,X_{t}^{\alpha})\Gamma(g_{t}(\cdot)\big))$ (3.25)
@@ 3. $(c)$
    $\alpha(\cdot)$ is $\mathbb{F}$-adapted and $\int^{t}_{0}\left|\alpha_{s}\right|^{2}ds<\infty$, $\forall\;t\in[0,T]$, $\mathbb{P}$-a.s..
@@ 3.1.2 The general correlation case: A verification argument
Since owing to Lemma 5.1, $\mathcal{E}\big(\int_{0}^{t}\alpha_{s}^{*\top}dB_{s}\big)$ is a true exponential martingale, we redefine the set of admissible portfolio strategy in this non-degenerate setting by
[…]
Remark: In particular, $\alpha\in{L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}$ and recalling that $\alpha_{t}:=\sigma(V_{t})^{\top}\pi_{t}=\sqrt{\operatorname{diag}(V_{t})}^{\top}\pi_{t}$ where $(\pi_{t})_{t\in[0,T]}$ is a deterministic $\mathbb{R}^{d}$-valued process representing a proportion of wealth, we can assume that $(\pi_{t})_{t\in[0,T]}$ is bounded by some constant $\pi^{\max}\in\mathbb{R}^{d}$ so that a sufficient condition for the exponential integrability in (3.53) is that Equation (3.31) admits a $C([0,T],\mathbb{R}^{d})$ solution for a sufficiently large constant $a\geq\|\pi^{\max}\|:=\|\pi\|_{\infty,T}$
We will directly show that the value attained by the portfolio strategy (3.52) is $\frac{x_{0}^{\gamma}}{\gamma}\Gamma_{0}$, and that no other admissible portfolio strategy can yield a larger value, through the following verification result:
@@ 1. $(a)$
    $\alpha(\cdot)$ is $\mathbb{F}$-adapted and $\mathbb{E}[\int^{t}_{0}\left|\alpha_{s}\right|^{2}ds]<\infty$, $\forall\;t\in[0,T]$;
@@ 3.2.1 The degenerate correlation case
$=\;\Big(\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\big(\theta_{i}+\rho\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\sqrt{V_{t}^{i}}\Big)_{1\leq i\leq d},\quad 0\leq t\leq T.$ (3.69)
@@
The proof is a straightforward adaption of the arguments from the proof of Theorem 3.12 to the particular case of degenerate correlation.
@@ 3.2.2 The general correlation case: A verification argument
First, note that if $K_{i}$ satisfies the Assumption 2.1 for $i=1,\ldots,d$. As $1-\rho_{i}^{2}\geq 0$, then (Gnabeyeu, 2026, Theorem A.2.) provides the existence of a unique global continuous solution on $[0,T]$ to (3.70)– (3.71). More details are given in the Remark on Proposition 3.11. In particular, when $\rho_{1}=\ldots=\rho_{d}=\rho$, the function $\psi$ coincides with the unique global solution to the Riccati–Volterra equations (3.62)–(3.63). Consequently, to avoid restrictions on the correlation structure, we use a verification approach for the Martingale optimality principle to solve the optimization problem, thus extending the results obtained in the preveous section to the more general correlation structure.
@@
For the sake of brevity, yet without compromising self-containment, we present a sketch of the proof, deferring the detailed proof to Section 5.
Sketch of Proof: We adopt a different verification approach from that of Section 3.1.2. In the spirit of Hu et al. (2005), we consider the family of stochastic processes $\{J^{\alpha}_{t}\}_{t\in[0,T]}$, $\alpha\in{\mathcal{A}}$ defined for every $t\in[0,T]$ by
$J_{t}^{\alpha}:=-\frac{1}{\gamma}\exp(-\gamma x_{0}e^{\int^{T}_{0}r(s)ds})\exp\!\Big(-\gamma\int_{0}^{t}e^{\int^{T}_{s}r(u)du}\big(\alpha_{s}^{\top}dB_{s}+\alpha_{s}^{\top}\lambda_{s}\mathrm{d}s\big)\Big)\Gamma_{t}=-\frac{\Gamma_{t}}{\gamma}\exp(-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t})$ (3.86)
where the pair $(\Gamma,\Lambda)$ satisfies the Riccati Backward SDE (BSDE) (3.74)–(3.75) under $\mathbb{P}$. We verify that $\{J^{\alpha}_{t}\}_{t\in[0,T]}$, $\alpha\in{\mathcal{A}}$ satisfies the Martingale optimality principle in the sense of Definition 3.2.